# Adaptive path planning for efficient object search by UAVs in agricultural fields

### v1

@@ 2.8.4 Experiment 4: Number of objects
To investigate the effect of different number of objects in the field on the flight path length, the number of objects in the field was altered. For each dataset, 140 additional objects were created by cropping annotated plants from the trainings dataset, to get to a total number of 200 objects. These cropped images were randomly rotated and added to the orthomosaic, creating realistic images for the detection network. Depending on the distribution in the original orthomosaic, the additional objects were distributed in the field using the following method:
@@
We tested the influence of the different number of objects by ranging the number of objects between 0 and 200 in steps of 20. Since each original orthomosaic contained 60 objects, we masked out object locations when we needed less than 60 objects by replacing the pixel values by a green color. When more than 60 objects were needed, we added them from the additional objects. For the clustered distribution, the object locations were sorted by cluster to make sure that we added or removed objects within a cluster first before adding or removing another cluster. Figure 7 shows an example of the original and added object locations on the dataset of 13 February 2024 for both the clustered and uniform distribution. For each number of objects, we compared the object density in number of objects per hectare with the relative difference in flight path between the adaptive planner and the baseline. This was done for the orthomosaics with a clustered and uniform distribution of objects separately.
@@ 3.1 Experiment 1: Relationship between detection certainty and errors
To investigate the ability of the different detection certainty measures to separate TP and FP detections, Figure 8 shows the distribution for the YOLOv8 confidence value and the certainty measures obtained using MCD for the TPs and FPs detections at different altitudes. A good certainty measure should result in clear distinction between TP detections with a high certainty and FP detections with a low certainty. The certainty for the TP detections was on average higher than for FP detections, except for the localization certainty at 24m and 32m altitude. For all certainty measures, an increasing altitude resulted in an increasing overlap between the certainty distributions of the TP and FP detections. The default YOLOv8 confidence value showed a clear distinction for TP and FP detections at 12m altitude, however, increasing the altitude increased the overlap between the TP and FP confidences. The MCD occurrence measure resulted in most cases in a certainty of 1.0, and therefore could not be used to separate TPs and FPs. The MCD class uncertainty measure showed a clear distinction between TPs and FPs at 12m altitude. The localization certainty showed a distinction between TP and FP detections for 12m altitude. However, the distributions at higher altitudes did show more overlap compared to the 12m altitude. The combined MCD certainty value did not show a better distinction between true and false positives than the default YOLOv8 confidence score.
@@
Table 2 shows the t-values resulting from Welch’s t-test for testing the alternative hypothesis that the mean confidence of true positives is higher than the mean confidence of the false positives. The higher the t-value, the better the certainty measure is in separating true and false positives. As could also be seen in Figure 8, an increasing altitude decreased separability between the TP and FP distribution, indicating a more difficult distinction between correct and wrong detections. The standard YOLOv8 certainty measure performed best on the 12m and 24m altitude. The MCD class performed best on the 32m altitude, with the default YOLOV8 confidence score as second-best.
@@
Using the MCD certainty measures did not resulted in a better separation between correct and wrong detections for 12m and 24m and the difference in performance between MCD class and the YOLOv8 confidence measure at 32m was small. Additionally, applying MCD increased the processing time by a factor 5. Therefore, the rest of the experiments uses the YOLOv8 confidence value as certainty measure. The average F1-scores for all images in the dataset, using the YOLOv8 confidence value as the certainty measure, were 0.83 at 12m, 0.70 at 24m, and 0.44 at 32m altitude.
Table 2: T-values from Welch’s t-test for standard yolo-v8 certainty and the certainty methods using Monte-Carlo Dropout (MCD), testing the alternative hypothesis that the mean confidence of true positives is higher than the mean confidence of the false positives. Values indicated with a ’*’ are significant ($\alpha=0.05$), the highest t-values for each altitude are indicated in bold.
@@ 3.2 Experiment 2: Parameter optimization
Figure 9 shows the influence of the coverage altitude ($h_{\textrm{cov}}$), acceptance confidence threshold ($c_{\textrm{accept}}$) and rejection confidence threshold ($c_{\textrm{reject}}$) on the F1-score for the adaptive planner and the coverage planner for both the clustered and the uniformly distributed plants. For the coverage planner, the F1-score decreased when the altitude increased. For the adaptive planner, the performance also decreased with an increasing altitude, however, to a lesser extend than the coverage planner depending on the rejection confidence threshold. Increasing the rejection confidence lowered the F1-score for both distributions, especially in combination with a high coverage altitude. Overall, the F1-scores were higher on the clustered distribution compared to the uniform distribution, however, the effect of changing the parameters was similar for both distributions.
@@
Figure 10 shows the influence of the coverage altitude, acceptance confidence threshold and the rejection confidence threshold on the relative path length (with respect to a coverage planner at 12m altitude) for the coverage and adaptive planner. Flying at 48m altitude decreases the path length for the coverage planner with around 70%. The adaptive planner yielded a longer flight path than the coverage planner at 48m altitude due to the addition of the inspection flight path, however, its still around 40% shorter than the 12m coverage flight path. Increasing the rejection confidence decreased the flight path length due to a lower number of objects that needed to be inspected at lower altitude. This effect was stronger for the uniform distribution, since the distance between the objects was on average larger than for the clustered distribution. A higher acceptance threshold resulted in a longer flight path, because more objects needed to be inspected at low-altitude. However, this effect was only visible at a coverage altitude of 24m.
@@
Table 3 shows the parameters that yielded both a good performance and having a relative short flight path length. These parameters were used in experiment 3 and 4. Additionally, it shows the corresponding F1-score for the adaptive planner, the coverage planner using the same coverage altitude, and the normalized flight path length. The coverage flight path has a higher F1-score than the adaptive flight path for both the clustered distribution and the uniform distribution, however, at cost of a longer flight path length. At cost of a 2% lower F1-score, the flight path length decreased with 37% for the clustered distribution. For the uniform distribution, the advantages of the adaptive path planner are smaller, at the cost of a 2% lower F1-score, the flight path length only decreased 6%. Figure 11 shows the resulting adaptive flight path for both the clustered and uniform distribution of objects on the four different recording dates.
Table 3: Selected parameters and the resulting F1-score and normalized flight path length for the clustered distribution and the uniform distribution having a high F1-score and a short flight path length.
@@
Figure 11: Flight paths for the adaptive planner using the optimized parameters corresponding to the datasets of 13 February 2024 (a, e), 23 April 2024 (b, f), 18 July 2024 (c, g) and 1 Augustus 2024 (d, h) with a clustered (a-d) and uniform (e-h) distribution of plants. The start and end location of the UAV are indicated with an arrow, the ground truth locations of weed type i and ii are indicated with red and blue markers respectively.
@@ 3.3 Experiment 3: Localization errors
Figure 12 shows the effect of localization errors on the performance of the adaptive planner and the low-altitude coverage planner for both the clustered and uniform distribution of objects. Five levels of error were simulated, from ’perfect’ to ’very poor’ (see section 2.8.3). A higher localization error resulted in a lower F1-score for both planners. However, since the absolute error in localization is smaller at a lower altitude, the adaptive path planner yielded a higher F1-score at 24m, 36m and 48m altitude than the coverage flight path as it used additional low-altitude inspections, thereby improving localization of the objects. There was a notable drop in performance between the levels ’decent’ and ’poor’, indicating a maximum allowable error of 0.03m in position and altitude, along with a 1-degree error in the UAV’s heading, gimbal roll, and pitch. With a higher localization error, the low F1-scores can be explained by a higher number of FPs due to multiple observations of the same plant not being fused together, resulting in multiple plants. There was no notable difference in the F1-score between both planners for the clustered and uniform distribution of plants.
@@ 3.4 Experiment 4: Number of objects
Figure 13 shows the effect of the number of objects in the field on the relative flight path length for the clustered and uniform distributed objects. For a clustered distribution of objects, the adaptive planner resulted in a shorter flight path than the low-altitude coverage planner for all tested number of objects. However, when there were more than 250 objects per hectare, the adaptive flight path length was almost similar. For the uniform distribution of plants, however, the adaptive planner yielded a longer flight path when there were more than 80 - 120 objects per hectare in the field compared to the low-altitude coverage planner. With the more uniform object distribution on this sized field, the adaptive planner has to inspect so many plants at low altitude, that is more efficient to use a full-coverage low-altitude flight path.
@@
Figure 14 shows the flight paths for the adaptive path planner and the coverage planner for the adapted dataset for the four different recording dates with 20, 80, 140 and 200 objects in the field. When the objects were uniformly distributed, the UAV had to visit locations which were more uniformly distributed over the field, requiring more waypoints and thereby increased the flight path length compared to the clustered distribution.
Figure 14: Flight paths for the adaptive planner corresponding to the datasets of 13 February 2024 with 20 objects (a, e), 23 April 2024 with 80 objects (b, f), 18 July 2024 with 140 objects (c, g) and 1 Augustus 2024 with 200 objects (d, h) with a clustered (a-d) and uniform (e-h) distribution of plants. The start and end location of the UAV are indicated with an arrow, the ground truth locations of weed type i and ii are indicated with red and blue markers respectively.
@@ 4 Discussion
The results showed that, especially when the objects are distributed in clusters in the field, there is a lot of potential in using an adaptive path planner compared to a coverage path planner. Using an adaptive path planner with the optimized parameters, as defined in Table 3, decreased the flight path length with 37% on our datasets while the accuracy only decreased 2% compared to the baseline when the objects were distributed in clusters. Compared to a coverage planner, the adaptive planner was more robust against localization errors (Figure 12). When increasing the number of objects in the field and the objects were distributed in clusters, the adaptive path planner outperformed a coverage path planner at 12m altitude till around 300 objects per hectare (Figure 13a). When the objects were uniformly distributed, however, the coverage planner outperformed the adaptive planner when there were more than 100 objects in the field (Figure 13b). In section 4.1 we discuss the detection certainty estimation, in section 4.2 the impact of the detection errors on the adaptive planner, section 4.3 the impact of the localization errors, section 4.4 the validity of the offline simulation method and finally, section 4.5 discusses the usability of the adaptive planner in practice.
@@ 4.1 Detection certainty estimation
Using Monte-Carlo Dropout for certainty estimation did not show a better separability between TPs and FPs in experiment 1, except at 32 meter altitude. In the context of active learning, applying certainty estimation using MCD has shown to estimate the most uncertain image for annotation (Blok et al., 2022; Sokolova et al., 2024), and has shown a positive correlation between certainty and detection accuracy. This could also be observed in experiment 1, where the TPs on average had a higher certainty than the FPs. In contrast to experiment 1, Myojin et al. (2019) observed a small increase in the F1-score when applying MCD uncertainty estimation to an adapted YOLOv3 network for lunar crater detection. However, the improvement was minimal, and since YOLOv8 is generally more accurate than YOLOv3 (Jegham et al., 2024), its confidence estimations are also more reliable.
A major disadvantage of sampling uncertainty using Monte-Carlo dropout is the added inference time. In our experiments, it increased the inference time with a factor of five, even with re-using intermediate results by the network. For the adaptive path planner, this could have impact on the efficiency, especially when using onboard processing with limited computational power. Using alternative methods for uncertainty calculation that are less computationally expensive, such as Laplacian approximation (Gawlikowski et al., 2023), might be considered for future work.
@@ 4.2 Impact of detection errors
Flying at a higher altitude increased the number of FN’s for the detection network and thereby resulted in a lower F1-score (Figure 9). When there is a FN, no low-altitude inspection could be planned, and therefore these objects were missed. Therefore, in experiment 2, the best working rejection confidence was the lowest of the tested values, 0.05, to have a minimum number of FNs. This resulted in more FP detections, however, they were corrected later on due to the low-altitude inspection. The impact of FN detections is stronger for the uniform distribution than for the clustered distribution, because for the clustered distribution, there will probably be a low-altitude inspection of some objects in the cluster. As long as some of the objects in the cluster are detected with a low detection certainty, most objects in the cluster will be detected by the low-altitude inspections, also the ones that were missed from the high altitude. Experiment 2 indicated this by a lower optimal coverage altitude for the uniform distribution which limited the number of FN’s during the high-altitude coverage path.

### v2

@@ 2.8.4 Experiment 4: Number of objects
To investigate the effect of different numbers of objects in the field on the flight path length, the number of objects in the field was altered. For each dataset, 140 additional objects were created by cropping annotated plants from the training dataset, to get a total of 200 objects. These cropped images were randomly rotated and added to the orthomosaic, creating realistic images for the detection network. Depending on the distribution in the original orthomosaic, the additional objects were distributed in the field using the following method:
@@
We tested the influence of the different number of objects by ranging the number of objects between 0 and 200 in steps of 20. Since each original orthomosaic contained 60 objects, we masked out object locations when we needed fewer than 60 objects by replacing the pixel values with a green color. When more than 60 objects were needed, we added them from the additional objects. For the clustered distribution, the object locations were sorted by cluster to make sure that we added or removed objects within a cluster first before adding or removing another cluster. Figure 7 shows an example of the original and added object locations on the dataset of 13 February 2024 for both the clustered and uniform distributions. For each number of objects, we compared the object density in the number of objects per hectare with the relative difference in flight path between the adaptive planner and the baseline. This was done for the orthomosaics with a clustered and uniform distribution of objects separately.
@@ 3.1 Experiment 1: Relationship between detection certainty and errors
To investigate the ability of the different detection certainty measures to separate TP and FP detections, Figure 8 shows the distribution for the YOLOv8 confidence value and the certainty measures obtained using MCD for the TP and FP detections at different altitudes. A good certainty measure should result in a clear distinction between TP detections with a high certainty and FP detections with a low certainty. The certainty for the TP detections was, on average, higher than for FP detections, except for the localization certainty at 24m and 32m altitude. For all certainty measures, an increasing altitude resulted in an increasing overlap between the certainty distributions of the TP and FP detections. The default YOLOv8 confidence value showed a clear distinction for TP and FP detections at 12m altitude; however, increasing the altitude increased the overlap between the TP and FP confidences. The MCD occurrence measure resulted in most cases in a certainty of 1.0, and therefore could not be used to separate TPs and FPs. The MCD class uncertainty measure showed a clear distinction between TPs and FPs at 12m altitude. The localization certainty showed a distinction between TP and FP detections for 12m altitude. However, the distributions at higher altitudes did show more overlap compared to the 12m altitude. The combined MCD certainty value did not show a better distinction between true and false positives than the default YOLOv8 confidence score.
@@
Table 2 shows the t-values resulting from Welch’s t-test for testing the alternative hypothesis that the mean confidence of true positives is higher than the mean confidence of the false positives. The higher the t-value, the better the certainty measure is in separating true and false positives. As could also be seen in Figure 8, an increasing altitude decreased separability between the TP and FP distributions, indicating a more difficult distinction between correct and wrong detections. The standard YOLOv8 certainty measure performed best at 12m and 24m altitude. The MCD class performed best at 32m altitude, with the default YOLOV8 confidence score as second-best.
@@
Using the MCD certainty measures did not result in a better separation between correct and wrong detections for 12m and 24m, and the difference in performance between the MCD class and the YOLOv8 confidence measure at 32m was small. Additionally, applying MCD increased the processing time by a factor of 5. Therefore, the rest of the experiments use the YOLOv8 confidence value as the certainty measure. The average F1-scores for all images in the dataset, using the YOLOv8 confidence value as the certainty measure, were 0.83 at 12m, 0.70 at 24m, and 0.44 at 32m altitude.
Table 2: T-values from Welch’s t-test for standard yolo-v8 certainty and the certainty methods using Monte-Carlo Dropout (MCD), testing the alternative hypothesis that the mean confidence of true positives is higher than the mean confidence of the false positives. Values indicated with a ’*’ show a significant ($\alpha=0.05$) difference between the true and false positive distributions; the highest t-values for each altitude are indicated in bold.
@@ 3.2 Experiment 2: Parameter optimization
Figure 9 shows the influence of the coverage altitude ($h_{\textrm{cov}}$), acceptance confidence threshold ($c_{\textrm{accept}}$) and rejection confidence threshold ($c_{\textrm{reject}}$) on the F1-score for the adaptive planner and the coverage planner for both the clustered and the uniformly distributed plants. For the coverage planner, the F1-score decreased when the altitude increased. For the adaptive planner, the performance also decreased with an increasing altitude, however, to a lesser extent than the coverage planner, depending on the rejection confidence threshold. Increasing the rejection confidence lowered the F1-score for both distributions, especially in combination with a high coverage altitude. Overall, the F1-scores were higher on the clustered distribution compared to the uniform distribution, however, the effect of changing the parameters was similar for both distributions.
@@
Figure 10 shows the influence of the coverage altitude, acceptance confidence threshold, and the rejection confidence threshold on the relative path length (with respect to a coverage planner at 12m altitude) for the coverage and adaptive planner. Flying at 48m altitude decreases the path length for the coverage planner by around 70%. The adaptive planner yielded a longer flight path than the coverage planner at 48m altitude due to the addition of the inspection flight path; however, it is still around 40% shorter than the 12m coverage flight path. Increasing the rejection confidence decreased the flight path length due to a lower number of objects that needed to be inspected at a lower altitude. This effect was stronger for the uniform distribution, since the distance between the objects was on average larger than for the clustered distribution. A higher acceptance threshold resulted in a longer flight path, because more objects needed to be inspected at low altitude. However, this effect was only visible at a coverage altitude of 24m.
@@
Table 3 shows the parameters that yielded both a good performance and a relatively short flight path length. These parameters were used in experiments 3 and 4. Additionally, it shows the corresponding F1-score for the adaptive planner, the coverage planner using the same coverage altitude, and the normalized flight path length. The coverage flight path has a higher F1-score than the adaptive flight path for both the clustered distribution and the uniform distribution, however, at the cost of a longer flight path length. At the cost of a 2% lower F1-score, the flight path length decreased by 37% for the clustered distribution. For the uniform distribution, the advantages of the adaptive path planner are smaller, at the cost of a 2% lower F1-score, the flight path length only decreased 6%. Figure 11 shows the resulting adaptive flight path for both the clustered and uniform distribution of objects on the four different recording dates.
Table 3: Selected parameters and the resulting F1-score and normalized flight path length for the clustered distribution and the uniform distribution, having a high F1-score and a short flight path length.
@@
Figure 11: Flight paths for the adaptive planner using the optimized parameters corresponding to the datasets of 13 February 2024 (a, e), 23 April 2024 (b, f), 18 July 2024 (c, g) and 1 Augustus 2024 (d, h) with a clustered (a-d) and uniform (e-h) distribution of plants. The start and end location of the UAV are indicated with an arrow; the ground truth locations of weed types I and II are indicated with red and blue markers, respectively.
@@ 3.3 Experiment 3: Localization errors
Figure 12 shows the effect of localization errors on the performance of the adaptive planner and the low-altitude coverage planner for both the clustered and uniform distribution of objects. Five levels of error were simulated, from ’perfect’ to ’very poor’ (see section 2.8.3). A higher localization error resulted in a lower F1-score for both planners. However, since the absolute error in localization is smaller at a lower altitude, the adaptive path planner yielded a higher F1-score at 24m, 36m, and 48m altitude than the coverage flight path, as it used additional low-altitude inspections, thereby improving the object localization. There was a notable drop in performance between the levels ’decent’ and ’poor’, indicating a maximum allowable error of 0.03m in position and altitude, along with a 1-degree error in the UAV’s heading, gimbal roll, and pitch. With a higher localization error, the low F1-scores can be explained by a higher number of FPs due to multiple observations of the same plant not being fused together, resulting in multiple plants. There was no notable difference in the F1-score between the clustered and uniform distribution of plants.
@@ 3.4 Experiment 4: Number of objects
Figure 13 shows the effect of the number of objects in the field on the relative flight path length for the clustered and uniformly distributed objects. For a clustered distribution of objects, the adaptive planner resulted in a shorter flight path than the low-altitude coverage planner for all tested numbers of objects. However, when there were more than 250 objects per hectare, the adaptive flight path length was almost similar. For the uniform distribution of plants, however, the adaptive planner yielded a longer flight path when there were more than 80 - 120 objects per hectare in the field compared to the low-altitude coverage planner. With a more uniform object distribution, the adaptive planner has to inspect so many plants at low altitude that it is more efficient to use a full-coverage low-altitude flight path.
@@
Figure 14 shows the flight paths for the adaptive path planner and the coverage planner for the adapted dataset for the four different recording dates with 20, 80, 140, and 200 objects in the field. When the objects were uniformly distributed, the UAV had to visit locations that were more uniformly distributed over the field, requiring more waypoints and thereby increasing the flight path length compared to the clustered distribution.
Figure 14: Flight paths for the adaptive planner corresponding to the datasets of 13 February 2024 with 20 objects (a, e), 23 April 2024 with 80 objects (b, f), 18 July 2024 with 140 objects (c, g) and 1 Augustus 2024 with 200 objects (d, h) with a clustered (a-d) and uniform (e-h) distribution of plants. The start and end location of the UAV are indicated with an arrow; the ground truth locations of weed types I and II are indicated with red and blue markers, respectively.
@@ 4 Discussion
The results showed that, especially when the objects are distributed in clusters in the field, there is a lot of potential in using an adaptive path planner compared to a coverage path planner. Using an adaptive path planner with the optimized parameters, as defined in Table 3, decreased the flight path length by 37% on our datasets while the accuracy only decreased 2% compared to the baseline when the objects were distributed in clusters. Compared to a coverage planner, the adaptive planner was more robust against localization errors (Figure 12). When increasing the number of objects in the field and the objects were distributed in clusters, the adaptive path planner outperformed a coverage path planner at 12m altitude till around 300 objects per hectare (Figure 13a). When the objects were uniformly distributed, however, the coverage planner outperformed the adaptive planner when there were more than 100 objects in the field (Figure 13b). In section 4.1 we discuss the results in context of existing literature, section 4.2 the detection certainty estimation, in section 4.3 the impact of the detection errors on the adaptive planner, section 4.4 the impact of the localization errors, section 4.5 the validity of the offline simulation method and finally, section 4.6 discusses the usability of the adaptive planner in practice.
4.1 Comparison with literature
The results in this paper were compared to a full coverage flight path, which is the most used approach in practice, where it yielded a 37% shorter flight path at the cost of a 2% lower F1-score. van Essen et al. (2024) showed a 74% shorter flight path for a learning-based adaptive path planner, at the cost of a 12% lower detection accuracy. Popović et al. (2017) showed that their path planner detects 85% of the objects in half the time of a coverage planner, while detecting all objects in approximately the same time as the coverage planner. They both show the trade-off between the flight time and the detection accuracy. Compared to those planners, our planner had a relatively small drop in performance, but also a smaller decrease in flight path length compared to a coverage flight. However, by selecting a different coverage altitude and confidence thresholds, it is possible to shift the balance towards a shorter flight path (experiment 2).
Using the adaptive path planner only had advantages when the objects were non-uniformly distributed. This was also confirmed by Popović et al. (2017) and van Essen et al. (2024), who both showed that the flight path was less efficient when the objects were uniformly distributed.
@@ 4.2 Detection certainty estimation
Using Monte-Carlo Dropout for certainty estimation did not show a better separability between TPs and FPs in experiment 1, except at 32 meter altitude. In the context of active learning, applying certainty estimation using MCD has been shown to estimate the most uncertain image for annotation (Blok et al., 2022; Sokolova et al., 2024), and has shown a positive correlation between certainty and detection accuracy. This could also be observed in experiment 1, where the TPs, on average, had a higher certainty than the FPs. In contrast to experiment 1, Myojin et al. (2019) observed a small increase in the F1-score when applying MCD uncertainty estimation to an adapted YOLOv3 network for lunar crater detection. However, the improvement was minimal, and since YOLOv8 is generally more accurate than YOLOv3 (Jegham et al., 2024), its confidence estimations are also more reliable.
A major disadvantage of sampling uncertainty using Monte-Carlo dropout is the added inference time. In our experiments, it increased the inference time by a factor of five, even with reusing intermediate results by the network. For the adaptive path planner, this could have an impact on the efficiency, especially when using onboard processing with limited computational power. Using alternative methods for uncertainty calculation that are less computationally expensive, such as Laplacian approximation (Gawlikowski et al., 2023), might be considered for future work.
@@ 4.3 Impact of detection errors
Flying at a higher altitude increased the number of FN’s for the detection network and thereby resulted in a lower F1-score (Figure 9). When there is an FN, no low-altitude inspection could be planned, and therefore, these objects were missed. Therefore, in experiment 2, the best working rejection confidence was the lowest of the tested values, 0.05, to have a minimum number of FNs. This resulted in more FP detections, however, they were corrected later on due to the low-altitude inspection. The impact of FN detections is stronger for the uniform distribution than for the clustered distribution, because for the clustered distribution, there will probably be a low-altitude inspection of some objects in the cluster. As long as some of the objects in the cluster are detected with a low detection certainty, most objects in the cluster will be detected by the low-altitude inspections, and also the ones that were missed from the high altitude. Experiment 2 indicated this by a lower optimal coverage altitude for the uniform distribution, which limited the number of FN’s during the high-altitude coverage path.